# Multi-Field Adaptive Retrieval

### v1

@@ Abstract
Document retrieval for tasks such as search and retrieval-augmented generation typically involves datasets that are unstructured: free-form text without explicit internal structure in each document. However, documents can have a structured form, consisting of fields such as an article title, message body, or HTML header. To address this gap, we introduce Multi-Field Adaptive Retrieval (mFAR), a flexible framework that accommodates any number of and any type of document indices on structured data. Our framework consists of two main steps: (1) the decomposition of an existing document into fields, each indexed independently through dense and lexical methods, and (2) learning a model which adaptively predicts the importance of a field by conditioning on the document query, allowing on-the-fly weighting of the most likely field(s). We find that our approach allows for the optimized use of dense versus lexical representations across field types, significantly improves in document ranking over a number of existing retrievers, and achieves state-of-the-art performance for multi-field structured data.
@@ 1 Introduction
The task of document retrieval has many traditional applications, like web search or question answering, but there has also been renewed spotlight on it as part of LLM workflows, like retrieval-augmented generation (RAG). One area of study is focused on increasing the complexity and naturalness of queries (Yang et al., 2018; Qi et al., 2019; Jeong et al., 2024; Lin et al., 2023). Another less studied area considers increased complexity of the documents (Jiang et al., 2024; Wu et al., 2024b). This represents a challenge forward from prior datasets for retrieval, like MS MARCO (Nguyen et al., 2016), which contain chunks of text that are highly related to the query. As a result, dense retrievers achieve success by embedding text into a vector representations (Karpukhin et al., 2020; Ni et al., 2022; Izacard et al., 2022) or searching over the documents via lexical match (Robertson et al., 1994). Relatedly, there are also prior methods (Gao et al., 2021; Chen et al., 2022) concerning the benefits of a hybrid representation, but these hybrid options are not mainstream solutions. We revisit both hybrid models and methods for retrieval of more complex documents.
@@
Our motivation for this direction derives from two observations: 1) documents do have structure: fields like titles, timestamps, headers, authors, etc. and queries can refer directly to this structure; and 2) a different scoring method may be beneficial for each of these fields, as not every field is necessary to answer each query. More specifically, our goal is to investigate retrieval on structured data. Existing work on retrieval for structured data with dense representations focus on directly embedding structured knowledge into the model through pretraining approaches (Li et al., 2023; Su et al., 2023), but we would like a method which can more flexibly combine existing pretrained models and scorers. Similarly, there has been prior interest in multi-field retrieval, although those focused on retrieval with lexical or sparse features or early neural models (Robertson et al., 1994; Zaragoza et al., 2004; Zamani et al., 2018).
@@
In this work, we demonstrate how multi-field documents can be represented through paired views on a per-field basis, with a learned mechanism that maps queries to weighted combinations of these views. Our method, Multi-Field Adaptive Retrieval (mFAR), is a retrieval approach that can accommodate any number of fields and any number of scorers (such as one lexical and one vector-based) for each field. Additionally, we introduce a lightweight component that adaptively weights the most likely fields, conditioned on the query. This allows us to overspecify the fields initially and let the model decide which scorer to use to score each field. mFAR obtains significant performance gains over existing state-of-the-art baselines. Unlike prior work, our simple approach does not require pretraining and offers some controllability at test-time. Concretely, our contributions are:
@@
    We introduce a novel framework for document retrieval, mFAR, that is aimed at structured data with any number of fields. Notably, mFAR is able to mix lexical and vector-based scorers between the query and the document’s fields.
@@
    We find that a hybrid mixture of scorers performs better than using dense or lexical-based scorers alone; we also find that encoding documents with our multi-field approach can result in better performance than encoding the entire document as a whole. As a result, mFAR achieves state-of-the-art performance on STaRK, a dataset for structured document retrieval.
@@
    Finally, we analyze the performance of our model trained from our framework; we control the availability of scorers at test-time in an ablation study to measure the importance of the individual fields in the corpus
@@
Figure 1: Traditional documents for retrieval (top), like in MS MARCO (Nguyen et al., 2016) and BioASQ (Nentidis et al., 2023), are unstructured: free-form text that tends to directly answer the queries. Documents in the STaRK datasets (bottom) (Wu et al., 2024b), are structured: each contains multiple fields. The queries require information from some of these fields, so it is important to both aggregate evidence across multiple fields while ignoring irrelevant ones.
@@ 2 Multi-field Retrieval
While structured documents is a broad term in general, in this work, we focus on documents that can be decomposed into fields, where each field has a name and a value. As an example in Figure 1, for the STaRK-Prime document, Entity Type would be a field name and its value would be “drug.” The values themselves can have additional nested structure, like Category has a list of terms as its value. Note that this formulation of structured multi-field document is broadly applicable, as it not only includes objects like knowledge base entries, but also free-form text (chat messages, emails) along with their associated metadata (timestamps, sender, etc).
Formally, we consider a corpus of documents $\mathcal{C}$ = {$d_{1}$, $d_{2}$, …, $d_{n}$} and a set of associated fields $\mathcal{F}$ = {$f_{1}$, $f_{2}$, …, $f_{m}$} that make up each document $d$, i.e., $d=\{f:x_{f}\mid f\in\mathcal{F}\}$, where $x_{f}$ is the value for that field. Then, given a natural-language query $q$, the goal is a scoring function $s(q,d)$ that can be used to rank the documents in $\mathcal{C}$ such that the relevant documents to $q$ are at the top (or within the top-$k$). $q$ may query about values from any subset of fields, either lexically or semantically.
@@ 2.1 Standard Retriever and Contrastive Loss
Traditionally, $d$ is indexed in its entirety. The retriever can employ either a lexical (Robertson et al., 1994) or dense (embedding-based) (Lee et al., 2019; Karpukhin et al., 2020) scorer. A lexical scorer like BM25 (Robertson et al., 1994) directly computes $s(q,d)$ based on term frequencies. For a dense scorer, document and query encoders are used to embed $d$ and $q$, and a simple similarity function, in our case an unnormalized dot product, is used to compute $s(q,d)$
The document and query encoders can be finetuned by using a contrastive loss (Izacard et al., 2022), which aims to separate a positive (relevant) document $d_{i}^{+}$ against $k$ negative (irrelevant) documents $d_{i}^{-}$ for a given query $q$. In prior work, a shared encoder for the documents and queries is trained using this loss, and a temperature $\tau$ is used for training stability:
$\mathcal{L}_{c}=-\log\frac{e^{s(q_{i},d_{i}^{+})/\tau}}{e^{s(q_{i},d_{i}^{+})/\tau}+\sum_{d_{i}^{-}}e^{s(q_{i},d_{i}^{-})/\tau}}$ (1)
$\mathcal{L}_{c}$ is the basic contrastive loss which maximizes $P(d_{i}^{+}\mid q_{i})$. Following Henderson et al. (2017) and Chen et al. (2020a), we employ in-batch negatives to efficiently sample those negative documents by treating the other positive documents $d_{j}^{+}$ where $j\neq i$ and $1\leq j\leq k$ as part of the negative set, $d_{i}^{-}$. Furthermore, following prior work (Yang et al., 2019; Ni et al., 2022), we can include a bi-directional loss for $P(q_{i}\mid d_{i}^{+})$. Here, for a given positive document $d_{j}^{+}$, $q_{j}$ is the positive query and the other queries $q_{i},i\neq j$ become negative queries:
$\mathcal{L}_{r}=-\log\frac{e^{s(q_{i},d_{i}^{+})/\tau}}{e^{s(q_{i},d_{i}^{+})/\tau}+\sum_{q_{j},j\neq i}e^{s(q_{j},d_{i}^{+})/\tau}}$ (2)
The final loss for the (shared) encoder is $\mathcal{L}=\mathcal{L}_{c}+\mathcal{L}_{r}$.
@@
Because structured documents can be decomposed into individual fields ($d=\{x_{f}\}_{f\in\mathcal{F}}$), we can score the query against each field separately. This score could be computed via lexical or dense (vector-based) methods. This motivates a modification to the standard setup above, where $s(q,d)$ can instead be determined as a weighted combination of field-wise scores and scoring methods,
@@
Here, $s_{f}^{m}(q,x_{f})$ is the score between $q$ and field $f$ of $d$ using scoring method $m$, and $\mathcal{M}$ is the set of scoring methods. For a hybrid model, =$\mathcal{M}=\{\text{lexical},\text{dense}\}$. $w_{f}^{m}$ is a weight, possibly learned, that is associated with field $f$ and scoring method $m$.
@@
As presented, our method uses weights, $w_{f}^{m}$, that are learned for each field and scorer. This is useful in practice, as not every field in the corpus is useful or even asked about, like ISBN numbers or internal identifiers. Additionally, queries usually ask about information contained in a small number of fields and these fields change depending on the query.
This motivates conditioning the value of $w_{f}^{m}$ also on $q$ so that the weights can adapt to the given query by using the query text to determine the most important fields. We define an adaptation function $G$ and let $w_{f}^{m}=G(q,f,m)$. Now, the query-conditioned, or adaptive, sum is:
@@
This can be implemented by learning embeddings for each field and scoring method, $\mathbf{a}_{f}^{m}$, and so $G(q,f,m)={\mathbf{a}_{f}^{m}}^{\top}\mathbf{q}$. We find that learning is more stable if we insert a nonlinearity: $G(q,f,m)=\text{softmax}(\{{\mathbf{a}_{f}^{m}}^{\top}\mathbf{q}\})$, and so this is the final version we use in mFAR.
@@
An objective of ours is to seamlessly incorporate scorers using different methods (lexical and dense). However, the distribution of possible scores per scorer can be on different scales. While $G$ can technically learn to normalize, we want $G$ to focus on query-conditioning. Instead, we experiment with using batch normalization (Ioffe & Szegedy, 2015) per field, that whitens the scores and learns new scalars $\gamma_{f}^{m}$ and $\beta_{f}^{m}$ for each field and scorer. Because these scores are ultimately used in the softmax of the contrastive loss, $\gamma_{f}^{m}$ acts like a bias term which modulates the importance of each score and $\beta_{f}^{m}$ has no effect.
Note that the score whitening process is not obviously beneficial or necessary, especially if the scorers do already share a similar distribution (i.e. if we only use dense scorers). We leave the inclusion of normalization as a hyperparameter included in our search.
@@
At test time, the goal is to rank documents by $s(q,d)$ such that the relevant (gold) documents are highest. Because it can be slow to compute $|\mathcal{F}||\mathcal{M}||\mathcal{C}|$ scores for the whole corpus, we use an approximation. We first determine a top-$k$ shortlist, $\mathcal{C}_{f}^{m}$, of documents for each field and scorer and only compute the full scores for all $\bigcup_{f\in\mathcal{F},m\in\mathcal{M}}\mathcal{C}_{f}^{m}$, which results in the final ranking. Note this inexact approximation of the top-$k$ document is distinct from traditional late-stage re-ranking methods that rescore the query with each document, which is not the focus of this work.
@@ 3 Experiments
Dataset Domain Num. Documents Num. Fields Train Dev. Test.
Amazon products, product reviews 950K 8 6K 1.5K 1.5K
MAG science papers, authors 700K 5 8K 2.6K 2.6K
Prime biomedical entities 130K 22 6.1K 2.2K 2.8K
@@ 3.1 Data
In the retrieval baselines (Wu et al., 2024b), node information is linearized into documents that can be encoded and retrieved via dense methods. We likewise treat each node as a document. In our work, we preserve each node property or relation as a distinct field for our multi-field models or likewise reformat to a human-readable document for our single-field models. Compared to Amazon and MAG, we notice that Prime contains a higher number of relation types, i.e. relatively more fields in Prime are derived from knowledge-graph relations than either Amazon or MAG, where document content is derived from a node’s properties. The corpus sizes and number of fields are listed in Table 1. For more details on dataset preprocessing and the exact the fields used for each dataset, see Appendix A.
We use trec_eval111 https://github.com/usnistgov/trec_eval. for evaluation and follow Wu et al. (2024b) by reporting Hit@$1$, Hit@$5$, Recall@$20$, and mean reciprocal rank (MRR). [^2] We omit Hit@$5$ in the main paper due to space (it is reported in Appendix C).
@@ 3.2 Baselines and Prior Work
We compare primarily to prior work on STaRK, which is a set of baselines established by Wu et al. (2024b) and more recent work by Wu et al. (2024a). Specifically, they include two vector similarity search methods that use OpenAI’s text-embedding-ada-002 model, ada-002 and multi-ada-002. Notably, the latter is also a multi-vector approach, although it only uses two vectors per document: one to capture node properties and one for relational information. We also include their two LLM-based re-ranking baselines on top of ada-002.333Note their re-ranking baselines are evaluated on a random 10% subset of the data due to cost. Although our work does not perform re-ranking, we add these results to show the superiority of finetuning smaller retrievers over using generalist large language models for reranking.
@@
Finally, we use a state-of-the-art the pretrained retrieval encoder, Contriever finetuned on MS MARCO (Izacard et al., 2022, facebook/contriever-msmarco) as a baseline for our dense scorer, which we subsequently continue finetuning on STaRK. We use BM25 (Robertson et al., 2004; Lù, 2024) as a lexical baseline. These use the single-field formatting described Section 3.1.

### v2

@@ Abstract
Document retrieval for tasks such as search and retrieval-augmented generation typically involves datasets that are unstructured: free-form text without explicit internal structure in each document. However, documents can have some structure, containing fields such as an article title, a message body, or an HTML header. To address this gap, we introduce Multi-Field Adaptive Retrieval (mFAR), a flexible framework that accommodates any number and any type of document indices on semi-structured data. Our framework consists of two main steps: (1) the decomposition of an existing document into fields, each indexed independently through dense and lexical methods, and (2) learning a model which adaptively predicts the importance of a field by conditioning on the document query, allowing on-the-fly weighting of the most likely field(s). We find that our approach allows for the optimized use of dense versus lexical representations across field types, significantly improves in document ranking over a number of existing retrievers, and achieves state-of-the-art performance for multi-field semi-structured data.
@@ 1 Introduction
The task of document retrieval has many traditional applications, like web search or question answering, but there has also been renewed interest as part of LLM workflows, like retrieval-augmented generation (RAG). An area of study is focused on increasing the complexity and naturalness of queries (Yang et al., 2018; Qi et al., 2019; Jeong et al., 2024; Lin et al., 2023). Another less studied area considers the increased complexity of the documents (Jiang et al., 2024; Wu et al., 2024b). This represents a challenge compared to prior datasets for retrieval, like MS MARCO (Nguyen et al., 2016), which contain chunks of text that are highly related to the query. Retrieval is done by either searching over the documents via lexical match (Robertson et al., 1994) or with dense retrievers that embed text into vector representations (Karpukhin et al., 2020; Ni et al., 2022; Izacard et al., 2022). Relatedly, some approaches (Gao et al., 2021; Chen et al., 2022) explore the benefits of a hybrid solution, but these options are not mainstream. In this work, we revisit both hybrid models and methods for retrieval of more complex documents.
@@
Our motivation for this direction derives from two observations: 1) documents do have structure: fields like titles, timestamps, headers, authors, etc. and queries can refer directly to this structure; and 2) a different scoring method may be beneficial for each of these fields, as not every field is necessary to answer each query. More specifically, our goal is to investigate retrieval on semi-structured data. Existing work on retrieval for semi-structured data with dense representations focus on directly embedding semi-structured knowledge into the model through pretraining approaches (Li et al., 2023; Su et al., 2024), but we would like a method which can more flexibly combine existing pretrained models and scorers. Similarly, there has been prior interest in multi-field retrieval, although these works focused on retrieval with solely lexical or sparse features or early neural models (Robertson et al., 1994; Zaragoza et al., 2004; Zamani et al., 2018).
@@
In this work, we demonstrate how multi-field documents can be represented through paired views and on a per-field basis, with a learned mechanism that maps queries to weighted combinations of these views. Our method, Multi-Field Adaptive Retrieval (mFAR),111 https://github.com/microsoft/multifield-adaptive-retrieval is a retrieval approach that can accommodate any number of fields and any number of scorers (such as one lexical and one vector-based) for each field. Additionally, we introduce a light-weight component that adaptively weights the most likely fields, conditioned on the query. This allows us to exhaustively include all fields and scorers at inference and let the model determine the relative importance. mFAR obtains significant performance gains over existing state-of-the-art baselines. Unlike prior work, our simple approach does not require pretraining and offers some controllability at test-time. Concretely, our contributions are:
@@
    We introduce a novel framework for document retrieval, mFAR, that is aimed at semi-structured data with any number of fields. Notably, mFAR is able to mix lexical and vector-based scorers between the query and the document’s fields.
@@
    We find that a hybrid mixture of scorers performs better than using dense or lexical-based scorers alone; we also find that encoding documents with our multi-field approach can result in better performance than encoding the entire document as a whole. As a result, mFAR achieves state-of-the-art performance on STaRK, a dataset for semi-structured document retrieval.
@@
    Finally, we analyze the performance of our models trained from our framework; we control the availability of scorers at test-time in an ablation study to measure the importance of the individual fields in the corpus.
@@
Figure 1: Traditional documents for retrieval (top), like in MS MARCO (Nguyen et al., 2016) and BioASQ (Nentidis et al., 2023), are unstructured: free-form text that tends to directly answer the queries. Documents in the STaRK datasets (bottom) (Wu et al., 2024b), are semi-structured: each contains multiple fields. The queries require information from some of these fields, so it is important to both aggregate evidence across multiple fields while ignoring irrelevant ones.
@@ 2 Multi-field Retrieval
While semi-structured documents is a broad term more generally, in this work, we focus on documents that can be decomposed into fields, where each field has a name and a value. As an example in Figure 1, for the STaRK-Prime document, Entity Type would be a field name and its value would be “drug.” The values themselves can have additional nested structure, like Category has a list of terms as its value. Note that this formulation of semi-structured multi-field document is broad, as it not only includes objects like knowledge base entries, but also free-form text (chat messages, emails) along with their associated metadata (timestamps, sender, etc) and tabular data.
Formally, we consider a corpus of documents $\mathcal{C}$ = {$d_{1}$, $d_{2}$, …, $d_{n}$} and a set of associated fields $\mathcal{F}$ = {$f_{1}$, $f_{2}$, …, $f_{m}$} that make up each document $d$, i.e., $d=\{f:x_{f}\mid f\in\mathcal{F}\}$, where $x_{f}$ is the value for that field. Then, given a natural-language query $q$, we would like a scoring function $s(q,d)$ that can be used to rank the documents in $\mathcal{C}$ such that the most relevant documents to $q$ score highest (or within the top-$k$). $q$ may ask about values from any subset of fields, either lexically or semantically.
@@ 2.1 Standard Retriever and Contrastive Loss
Traditionally, $d$ is indexed in its entirety. The retriever can employ either a lexical (Robertson et al., 1994) or dense (embedding-based) (Lee et al., 2019; Karpukhin et al., 2020) scorer. A lexical scorer like BM25 (Robertson et al., 1994) directly computes $s(q,d)$ based on term frequencies. For a dense scorer, document and query encoders are used to embed $d$ and $q$, and a simple similarity function, in our case an unnormalized dot product, is used to compute $s(q,d)$.
Document and query encoders can be finetuned by using a contrastive loss (Izacard et al., 2022), which aims to separate a positive (relevant) document $d_{i}^{+}$ against $k$ negative (irrelevant) documents $\mathcal{D}_{i}^{-}$ for a given query $q$. In prior work, a shared encoder for the documents and queries is trained using this loss, and a temperature $\tau$ is used for training stability:
$\mathcal{L}_{c}=-\log\frac{e^{s(q_{i},d_{i}^{+})/\tau}}{e^{s(q_{i},d_{i}^{+})/\tau}+{\displaystyle\sum}_{d_{i}^{-}\in\mathcal{D}_{i}^{-}}e^{s(q_{i},d_{i}^{-})/\tau}}$ (1)
$\mathcal{L}_{c}$ is the basic contrastive loss which maximizes $P(d_{i}^{+}\mid q_{i})$. Following Henderson et al. (2017); Izacard et al. (2022) and Chen et al. (2020a), we employ in-batch negatives to efficiently sample those negative documents by treating the other positive documents in the batch (of $b$ documents), $d_{j}^{+}$, as negatives and including them into $\mathcal{D}_{i}^{-}$, where, $j\neq i$ and $1\leq j\leq b$. Furthermore, following prior work (Yang et al., 2019; Ni et al., 2022; Chen et al., 2025), we can include a bi-directional loss for $P(q_{i}\mid d_{i}^{+})$. Here, for a given positive document $d_{j}^{+}$, $q_{j}$ is the positive query and the other queries $q_{i},i\neq j$ become negative queries:
$\mathcal{L}_{b}=-\log\frac{e^{s(q_{i},d_{i}^{+})/\tau}}{e^{s(q_{i},d_{i}^{+})/\tau}+{\displaystyle\sum}_{q_{j},j\neq i}e^{s(q_{j},d_{i}^{+})/\tau}}$ (2)
The final loss for the (shared) encoder is $\mathcal{L}=\mathcal{L}_{c}+\mathcal{L}_{b}$.
@@
Because semi-structured documents can be decomposed into individual fields ($d=\{x_{f}\}_{f\in\mathcal{F}}$), we can score the query $q$ against each field separately. This score could be computed via lexical or dense (vector-based) methods. This motivates a modification to the standard setup above, where $s(q,d)$ can instead be determined as a weighted combination of field-wise scores and scoring methods,
@@
Here, $s_{f}^{m}(q,x_{f})$ is the score between $q$ and field $f$ of $d$ using scoring method $m\in\mathcal{M}$, and $\mathcal{M}$ is the set of scoring methods. For a hybrid model, =$\mathcal{M}=\{\text{lexical},\text{dense}\}$. $w_{f}^{m}$ is a weight, possibly learned, that is associated with field $f$ and scoring method $m$.
@@
As presented, our method uses weights, $w_{f}^{m}$, that are learned for each field and scorer. This is useful in practice, as not every field in the corpus is useful or even asked about, like unrelated numbers or internal identifiers. Additionally, queries usually ask about information contained in a small number of fields and these fields change depending on the query.
This motivates conditioning the value of $w_{f}^{m}$ also on $q$ so that the weights can adapt to the given query by using the query text to determine the most important fields. We use an adaptation function $G$ and let $w_{f}^{m}=G(q,f,m)$. Now, the query-conditioned, or adaptive, weighted sum is:
@@
To implement $G$, let $\mathbf{q}$ be a dense embedding of $q$, and $\mathbf{a}_{f}^{m}\in\mathbb{R}^{|\mathbf{q}|}$ be learnable parameters. Then we could define $G(q,f,m)={\mathbf{a}_{f}^{m}}^{\top}\mathbf{q}$. We find that learning is more stable with a nonlinearity over all fields $f$ and scorers $m$: $G(q,f,m)=\text{softmax}(\{{\mathbf{a}_{f}^{m}}^{\top}\mathbf{q}\})$, which is what we use in mFAR.
@@
One objective of ours is to seamlessly incorporate scorers using different methods (lexical and dense). However, the distribution of possible scores per scorer can be on different scales. While $G$ can technically learn to normalize, we want $G$ to focus on query-conditioning. Instead, we experiment with using batch normalization (Ioffe & Szegedy, 2015) per field that whitens the scores and learns new scalars $\gamma_{f}^{m}$ and $\beta_{f}^{m}$ for each field and scorer. Because these scores are ultimately used in the softmax of the contrastive loss, $\gamma_{f}^{m}$ acts like a bias term which modulates the importance of each score while $\beta_{f}^{m}$ has no effect.
Note that the score whitening process is not obviously beneficial or necessary, especially if the scorers already share a similar distribution (i.e. if we only use dense scorers). We leave the inclusion of normalization as a hyperparameter as part of our grid search.
@@
At test time, the goal is to rank documents by $s(q,d)$ such that the relevant (gold) documents are highest. Because it can be slow to compute $|\mathcal{F}||\mathcal{M}||\mathcal{C}|$ scores for the whole corpus, we use an approximation. We first determine a top-$k$ shortlist, $\mathcal{C}_{f}^{m}$, of documents for each field and scorer and only compute the full scores for all ${\displaystyle\bigcup}_{f\in\mathcal{F},m\in\mathcal{M}}~{}\mathcal{C}_{f}^{m}$, which results in the final ranking. Note this inexact approximation of the top-$k$ document is distinct from traditional late-stage re-ranking methods that rescore the query with each document, which is not the focus of this work.
@@ 3.1 Data
In the retrieval baselines (Wu et al., 2024b), node (corresponding to an entity in the knowledge graph) information is linearized into documents that can be encoded and retrieved via dense methods. We likewise treat each node as a document. In our work, we preserve each node property or relation as a distinct field for our multi-field models or likewise reformat to a human-readable document for our single-field models. Compared to Amazon and MAG, we notice that Prime contains a higher number of relation types, i.e. relatively more fields in Prime are derived from knowledge-graph relations than in either Amazon or MAG, where document content is derived from a node’s properties. In total, there are 22, 8, and 5 fields for Prime, Amazon, and MAG respectively; more details on dataset sizes, preprocessing, exact fields are described in Appendix 5.
We use trec_eval222 https://github.com/usnistgov/trec_eval. for evaluation and follow Wu et al. (2024b) by reporting Hit@$1$, Recall@$20$, and mean reciprocal rank (MRR). Hit@5 is reported in Appendix C.1 due to space limitations here.
@@ 3.2 Baselines and Prior Work
We compare primarily to prior work on STaRK, which is a set of baselines established by Wu et al. (2024b) and more recent work by Wu et al. (2024a). Specifically, they include two vector similarity search methods that use OpenAI’s text-embedding-ada-002 model, ada-002 and multi-ada-002. Notably, the latter is also a multi-vector approach, although it only uses two vectors per document: one to capture node properties and one for relational information. We also include their two LLM-based re-ranking baselines (Claude3 and GPT4 rerankers) on top of ada-002. Although our work does not perform re-ranking, we add these results to show the superiority of finetuning smaller retrievers over using generalist LLMs for reranking.
@@
Finally, we use an off-the-shelf pretrained retrieval encoder, Contriever finetuned on MS MARCO333 https://huggingface.co/facebook/contriever-msmarco.(Izacard et al., 2022), as a baseline for our dense scorer, which we subsequently continue finetuning on STaRK. Early experiments showed that Contriever performed better than other dense retrievers. We use BM25 (Robertson et al., 1994; Lù, 2024) as a lexical baseline. These use the single-field formatting described Section 3.1.